OpenStudy (anonymous):

solve the equation using quadratic formula. 2x (sqaured) +13x=0

OpenStudy (anonymous):

can some one please help ,e i dont get quadratic functions

pooja195 (pooja195):

Find a GCF first do you see any?

OpenStudy (anonymous):

13?

OpenStudy (abdullahm):

It says to solve by using the Quadratic Formula... What is the Quadratic Formula?

OpenStudy (anonymous):

"ax2 + bx + c = 0"

pooja195 (pooja195):

Thats standered form

OpenStudy (abdullahm):



OpenStudy (anonymous):

yeah but for this question thers no c

pooja195 (pooja195):

so that would be 0 then :)

OpenStudy (anonymous):

but dont u have tp plug in the numbers? or is it when theres no c then the answer is equivalent to 0?

pooja195 (pooja195):

Name your abc values

OpenStudy (anonymous):

a=2 b=13

pooja195 (pooja195):

\[\huge~\rm~ax^2+bx+c=0\] \[\huge~\rm~2x^2+13x+0=0\]

pooja195 (pooja195):

Now name the abc values

OpenStudy (anonymous):

a=2 b=13 c=0?

OpenStudy (abdullahm):

Great job so far :)

pooja195 (pooja195):

yes! :) \[\huge~\rm~~x=\frac{ -(13)~\pm \sqrt{(13)^2-4(2)(0)} }{ 2(2)}\]

OpenStudy (anonymous):

\[so.... x= -13 \pm \sqrt{169}\]

OpenStudy (anonymous):

and then i would subtract 8?

OpenStudy (anonymous):

i dont understnd now...

pooja195 (pooja195):

\[\huge~\rm~~x=\frac{ -13~\pm \sqrt{169~x~1} }{ 4}\] \[\huge~\rm~~x=\frac{ -13~\pm \sqrt{169} }{ 4}\] Whats the square root of 169?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

right?

pooja195 (pooja195):

\[\huge~\rm~~x=\frac{ -13~\pm~13 \sqrt{1} }{ 4}\] now set up two equations and solve \[\huge~\rm~~x=\frac{ -13~+~13 \sqrt{1} }{ 4}\] \[\huge~\rm~~x=\frac{ -13~-~13 \sqrt{1} }{ 4}\]

OpenStudy (anonymous):

so it would be -13/2 and 0?

OpenStudy (anonymous):

those are my two answers correct?

OpenStudy (anonymous):

so would i do the same thing for 6x (sqaured) +2x=4?

pooja195 (pooja195):

You would subtract the 4 \[\huge~\rm~6x^2+2x-4=0\] and go from that